What transition in the hydrogen spectrum would have the same wavelength as the Balmer transition from n=4 of He+ spectrum. [Rh=109700cm^-1]
1
Expert's answer
2013-07-26T08:10:52-0400
For Hydrogen Z=1, so we have that transitional energy is proportional to 1/1^2 - 1/2^2=3/4; for Helium Z=2, so the transitional energy is proportional to 2^2(1/2^2 - 1/4^2)=3/4. Hence, we get that the transition from n=2 to n=1 will have the same wavelength as energy which is inversely to wavelength.